1.What is the mass in grams of three atoms of sodium?  23Na. Molar mass = A in g/mol


1
Expert's answer
2021-12-30T08:48:04-0500

n = m / M

m = n x M

M (Na) = 23 g/mol

1 mole of matter contains 6.022 × 1023 elementary entities of the substance.

n (Na) = (1 x 3) / 6.022×1023 = 0.5 x 10-23 mol

m (Na) = 0.5 x 10-23 x 23 = 11.5 x 10-23 g